OpenStudy (anonymous):

what is the radius if the circumference is 13.4?

OpenStudy (anonymous):

Circumference = 2*pi*r = 13.4

OpenStudy (anonymous):

U still having trouble?

OpenStudy (anonymous):

r= 13.4 / (2*pi)